Calculate the distance of the point in front of the cornea at which parallel light originating inside the reduced eye is focused


1
Expert's answer
2021-06-22T09:48:03-0400

"\frac 1F=\frac{1}{d_0}-\frac 1f,"

"d_0=25~cm,"

"F=\infin,"

"\frac{1}{d_0}=\frac 1f,"

"f=d_0=25~cm."